Circular No. 8287

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004T, 2004U, 2004V, AND 2004X Further to IAUC 8285, H. Pugh and W. Li report the LOSS discov-

ery, on a KAIT image taken on Feb. 12.4 UT, of an apparent supernova (mag ∼ 18.1), designated SN 2004X and located at α = 11^{h}19^{m}15^{s}.05, δ = −15^{o}44′04′′.7 (equinox 2000.0), which is 2′′.9 west and 2′′.3 south of the nucleus of the apparent host galaxy. A KAIT image taken on 2003 June 6.2 showed nothing at this position (limiting mag ∼ 19.0). A. V. Filippenko, R. Chornock, and R. J. Foley, University of California, Berkeley, report that inspection of CCD spectra (range 310–890 nm), obtained on Feb. 13 with the Keck I 10-m telescope (+ LRIS), shows that SN 2004X is of type II, with prominent P-Cyg profiles typical of the type-II-plateau variety.

Filippenko, Chornock, and Foley report that inspection of Keck I CCD spectra (range 310–920 nm), obtained on Feb. 13 UT, shows that SN 2004V (IAUC 8285) is of type II, probably close to maximum brightness, with a very blue and nearly featureless continuum. Weak Hα emission (FWHM = 2700 km/s) is present, while other weak lines (e.g., He I 587.6 nm, Hβ, Hγ) have P-Cyg profiles. The recession velocity of the host galaxy, measured from a fit to the broad Hα emission line, is ≈ 12500 km/s.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004T (cf. IAUC 8285), obtained by M. Calkins on Feb. 13.43 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova. The spectrum consists of a blue continuum with P-Cyg lines of hydrogen and helium. Adopting the NASA/IPAC Extragalactic Database recession velocity of 6437 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 7600 km/s. A spectrum of SN 2004U (cf. IAUC 8285), obtained on Feb. 13.44, shows it to be a type-II supernova; the spectrum consists of a blue continuum with weak Balmer lines. A spectrum of SN 2004V, obtained by Calkins on Feb. 13.46, shows it to be a type-II supernova. The spectrum consists of a very blue continuum with incipient P-Cyg lines of hydrogen.

IRAS 05436–0007 K. Hornoch, Lelekovice, Czech Republic, reports that co-added R-band

CCD images taken of IRAS 05436–0007 (cf. IAUC 8284) on Feb. 11.913 UT show that the nebula extends 65′′ in α and 80′′ in δ, with a total mag of R = 13.21 ± 0.05 (measured in a circular aperture of diameter 80′′).

2004 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green